OpenStudy (anonymous):

the spreadsheet shows the monthly revenue and expenses for a new business.use the models to predictt the month in which revenue will equal expense. month revenue expenses jan 4000 22,000 feb 9000 24,000 mar 13,000 25,000 apr 16,000 27,000 may 21,000 30,000 Would someone please help and explain how to figure this out

OpenStudy (anonymous):

r=4100x-300 e=1900x-19900 oct. 4100x=300 1900X+19900 sept. r=4100x+300 e=1900x+19900 aug. r=4100x-300 e=1900+19900 sept. Not sure how to figure the one that is correct,please help